# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ A.3 Final Lemmas
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Furthermore, in this case at least two vertices of $\{u,v,w\}$ are in a 2EC component in $\mathsf{OPT}_{2}$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{vz,zy\}$, one of the dummy edges $\{uy,wy\}$, and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
@@
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{opt}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$.
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
Let $D^{\mathsf{C3}}\coloneq\{uv,uv,vw,vw\}$ denote the set of dummy edges used for the construction of $G_{2}^{\mathsf{C3}}$. First observe that the definition of Algorithm 2, Lemma 33, and the induction hypothesis gives
@@
We distinguish the following cases depending on which type combination of Proposition 18 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
@@
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Lemma 22 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C3}}+3$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with at most four dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm 2 did not apply, we conclude that $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C3}}+2$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with at most three dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{C2},\{\mathsf{A},\mathsf{B1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+1$ because we can turn $\mathsf{OPT}_{2}$ with at most one dummy edge of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives

### v2

@@ A.2.2 Auxiliary Lemmas for Algorithm˜3
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm˜3 and Definition˜16 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Furthermore, in this case at least two vertices of $\{u,v,w\}$ are in a 2EC component in $\mathsf{OPT}_{2}$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{vz,zy\}$, one of the dummy edges $\{uy,wy\}$, and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
@@
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{opt}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm˜3 and Definition˜16 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$.
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})+\alpha$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})+\alpha$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})+\alpha$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
@@
Let $D^{\mathsf{C3}}\coloneq\{uv,uv,vw,vw\}$ denote the set of dummy edges used for the construction of $G_{2}^{\mathsf{C3}}$. First observe that the definition of Algorithm˜3, Lemma˜44, and the induction hypothesis gives
@@
We distinguish the following cases depending on which type combination of Proposition˜17 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
@@
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Corollary˜34 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C3}}+3$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with at most four dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm˜3 did not apply, we conclude that $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C3}}+2$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with at most three dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm˜3 and Definition˜16 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm˜3 and Definition˜16 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
    $(\mathsf{C2},\{\mathsf{A},\mathsf{B1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm˜3 and Definition˜16 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+1$ because we can turn $\mathsf{OPT}_{2}$ with at most one dummy edge of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
@@
A.3 Proof of Definition˜5
In this final subsection, we prove Definition˜5. The statement is implied by the following three lemmas. We first show that $\mathsf{RED}(G)$ runs in polynomial time. The proof is similar to the corresponding proof in [GGA23].
Lemma 46 (running time).
Proof.
Let $n=|V(G)|$ and $m=|V(G)|$, and $\sigma=n^{2}+m^{2}$ be the size of the problem. We first argue that every non-recursive step of Algorithms˜1, 2 and 3 can be performed in time polynomial in $\sigma$. Clearly, all computations performed in Algorithm˜1 can be performed in polynomial time in $\sigma$. In particular, it is possible to check in polynomial-time whether $G$ contains an $\alpha$-contractible subgraph with at most $\frac{4}{\varepsilon}$ vertices: Enumerate over all subsets $W$ of vertices of the desired size; for each such $W$, one can compute in polynomial time a minimum-size 2-ECSS $\mathsf{OPT}^{\prime}$ of $G[W]$, if any such subgraph exists. Furthermore, one can compute in polynomial time a maximum cardinality subset of edges $F$ with endpoints in $W$ such that $G\setminus F$ is 2EC. Then, it is sufficient to compare the size of $\mathsf{OPT}^{\prime}$, if exists, with $|E(G[W])\setminus F|$.
We now consider Algorithm˜3. We already argued that the computation in Algorithm˜3 can be executed in polynomial time in $\sigma$. Moreover, determining the $\mathsf{C2}$ subcases in Algorithms˜3, 3 and 3 and selecting a desired $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution in Algorithms˜3 and 3 and $\mathsf{OPT}_{1}^{\mathsf{C3}}$ solution in Algorithm˜3 can be done in constant time, as $G_{1}$ is of constant size in this case. Finding the corresponding sets of edges $F$ in Algorithm˜3 can also be done in polynomial time in $\sigma$ by enumeration, because Lemmas˜37, 38, 39, 40, 41, 42, 43 and 44 guarantee that there exists a desired set of edges $F$ of constant size. The same arguments can also be used to argue that Algorithm˜2 can be implemented in polynomial time.
To bound the total running time including recursive steps, let $T(\sigma)$ be the running time of Algorithm˜1 as a function of $\sigma$. In each step, starting with a problem of size $\sigma$, we generate either a single subproblem of size $\sigma^{\prime}<\sigma$ (to easily see this in Algorithms˜2 and 3, recall that $2\leq|V_{1}|\leq|V_{2}|$ and $7\leq|V_{1}|\leq|V_{2}|$, respectively) or two subproblems of size $\sigma^{\prime}$ and $\sigma^{\prime\prime}$ such that $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Thus, $T(\sigma)\leq\max\{T(\sigma^{\prime}),T(\sigma^{\prime})+T(\sigma^{\prime\prime})\}+\mathrm{poly}(\sigma)$ with $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Therefore, $T(\sigma)\leq\sigma\cdot\mathrm{poly}(\sigma)$. ∎
Lemma 47 (correctness).
Proof.
[…]
If the condition in Algorithm˜1 applies, let $G_{i}=G[V_{i}\cup\{v\}]$. Note that both $G_{1}$ and $G_{2}$, which must be 2EC since $G$ is 2EC, each contain at most |V(G)|1|V(G)|-1 vertices, hence by induction hypothesis $\mathsf{RED}(G_{1})$ and $\mathsf{RED}(G_{2})$ are 2EC spanning subgraphs of $V_{1}\cup\{v\}$ and $V_{2}\cup\{v\}$, respectively. Thus, their union is a 2EC spanning subgraph of $G$. If the conditions in Algorithm˜1 or Algorithm˜1 apply, $G^{\prime}\coloneq G\setminus\{e\}$ must be 2EC by Lemmas˜19 and 20, respectively, and $|V(G^{\prime})|=|V(G)|$ and $|E(G^{\prime})|=|E(G)|-1$. Hence, by induction hypothesis $\mathsf{RED}(G^{\prime})$ is a 2EC spanning subgraph of $G^{\prime}$, and thus also for $G$. If the condition in Algorithm˜1 applies, the claim follows from ˜16.
It remains to consider the cases when Algorithm˜2 or Algorithm˜3 is executed. Observe that in these cases, $G$ is 2VC since the condition of Algorithm˜1 is checked before. Moreover, note that all recursive calls to $\mathsf{RED}$ made in Algorithm˜2 and Algorithm˜3 consider graphs with strictly fewer vertices than $G$, hence the induction hypothesis applies. If Algorithm˜1 executes Algorithm˜2, the correctness follows from the induction hypothesis, Corollary˜25 and Lemmas˜27, 28 and 29. If Algorithm˜1 executes Algorithm˜3, the correctness follows from the induction hypothesis, Corollary˜25 and Lemmas˜38, 39, 40, 41, 42, 43 and 44. ∎
The following lemma implies $\mathsf{red}(G)\leq(\alpha+4\varepsilon)\mathsf{opt}(G)$ using the trivial bound $\mathsf{opt}(G)\geq|V(G)|$.
Lemma 48 (approximation guarantee).
For any 2-ECSS instance $G$ it holds that
Proof.
[…]
If the condition in Algorithm˜1 applies, recall that by Lemma˜18 we have $\mathsf{opt}(G)=\mathsf{opt}(G_{1})+\mathsf{opt}(G_{2})$, where $G_{i}\coloneq G[V_{i}\cup\{v\}]$ for $i\in\{1,2\}$. We distinguish a few cases depending on the sizes of $V_{1}$ and $V_{2}$. Assume w.l.o.g. $|V_{1}|\leq|V_{2}|$. If $|V_{2}|\leq\frac{4}{\varepsilon}-1$, then $\mathsf{red}(G)=\mathsf{opt}(G_{1})+\mathsf{opt}(G_{2})=\mathsf{opt}(G)\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon|V(G)|-4$ since $\alpha\geq 1$ and $|V(G)|\geq\frac{4}{\varepsilon}$. Otherwise, $|V_{2}|\geq\frac{4}{\varepsilon}$. If $|V_{1}|\leq\frac{4}{\varepsilon}-1$, then by the induction hypothesis we have $\mathsf{red}(G)\leq\mathsf{opt}(G_{1})+\alpha\mathsf{opt}(G_{2})+4\varepsilon|V(G_{2})|-4\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4$ using $\alpha\geq 1$ and $|V(G_{2})|\leq|V(G)|$. Otherwise, $|V_{2}|\geq|V_{1}|\geq\frac{4}{\varepsilon}$, and the induction hypothesis gives $\mathsf{red}(G)\leq\alpha\mathsf{opt}(G_{1})+\alpha\mathsf{opt}(G_{2})+4\varepsilon(|V(G_{1})|+|V(G_{2})|)-4\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|+4\varepsilon-8$, which implies the stated bound as $4\varepsilon\leq 4$.
If the condition in Algorithm˜1 applies, we have $\mathsf{opt}(G)=|\mathsf{OPT}(G)\cap{V(H)}^{2}|+|\mathsf{OPT}(G)\setminus{V(H)}^{2}|\geq\frac{1}{\alpha}|H|+\mathsf{opt}(G|H)$ by the definition of a contractible subgraph $H$ and observing that $\mathsf{OPT}(G)\setminus{V(H)}^{2}$ induces a feasible 2EC spanning subgraph of $G|H$. If $|V(G|H)|\leq\frac{4}{\varepsilon}$ one has $\mathsf{red}(G)=|H|+\mathsf{opt}(G|H)\leq\alpha\mathsf{opt}(G)\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4$ since $|V(G)|\geq\frac{4}{\varepsilon}$ whenever Algorithm˜1 reaches Algorithm˜1. Otherwise, we conclude that $\mathsf{red}(G)\leq|H|+\alpha\mathsf{opt}(G|H)+4\varepsilon|V(G|H)|-4\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4$ since $|V(G)|\geq|V(G|H)|$.
Finally, if Algorithm˜1 executes Algorithm˜2 or Algorithm˜3, we have that $|V(G)|>\frac{4}{\varepsilon}$. Thus, in these cases the statement follows from the induction hypothesis and Lemma˜30 or Lemma˜45, respectively. ∎